# A Unified Bilevel Model for Adversarial Learning and A Case Study

### v1

@@ Abstract
Adversarial learning has been attracting more and more attention thanks to the fast development of machine learning and artificial intelligence. However, due to the complicated structure of most machine learning models, the mechanism of adversarial attacks is not well interpreted. How to measure the effect of attack is still not quite clear. In this paper, we propose a unified bilevel model for adversarial learning. We further investigate the adversarial attack in clustering models and interpret it from data perturbation point of view. We reveal that when the data perturbation is relatively small, the clustering model is robust, whereas if it is relatively large, the clustering result changes, which leads to an attack. To measure the effect of attacks for clustering models, we analyse the well-definedness of the so-called $\delta$-measure, which can be used in the proposed bilevel model for adversarial learning of clustering models.
@@ 1 Introduction
With the fast development of machine learning and artificial intelligence, adversarial learning is receiving growing attentions. The first striking demonstration came from Szegedy et al. [1], who showed that imperceptible perturbations can reliably force modern neural networks to misclassify inputs, thereby exposing a fundamental vulnerability of high-capacity models. Goodfellow et al. [2] provided a concise explanation and practical attack algorithm so called the Fast Gradient Sign Method (FGSM) for fast, effective adversarial example synthesis, and thus enabled a set of adversarial-training defenses. Kurakin et al. [3] showed that adversarial training can be implemented on larger-scale datasets such as ImageNet, and further revealed that this approach leads to a significant improvement in the robustness of one-step methods. Roberts and Smyth [4] analyzed stochastic gradient descent under Byzantine adversaries, highlighting robustness issues that also arise in distributed or large-scale settings. Chhabra et al. [5] presented a black-box adversarial attack algorithm for clustering models with linearly separable clusters. Later Chhabra et al. [6] proposed a black-box adversarial attack against deep clustering models . We refer to [7, 8, 9] for the review and monographs of adversarial learning.
@@
Following widely adopted taxonomies, adversarial attacks can be divided into three types based on attacker’s knowledge: white-box (full access to parameters/gradients), black-box (only queries or input–output pairs), and gray-box in between. Due to different goals of attacks, it can also be classified as confidence reduction, untargeted misclassification and targeted misclassification. Madry et al. [10] casted adversarial training as min–max robust optimization and used multi-step PGD as a strong white-box baseline for $l_{p}$-bounded threats. Dong et al. [11] added momentum to iterative gradients to markedly improve black-box transferability over standard iterative attacks. Brendel et al. [12] proposed boundary attack that needs only top-1 decisions, walking along the boundary while shrinking perturbations. Papernot et al. [13] targeted a specific label by perturbing a small, saliency chosen set of pixels (sparse $L_{0}$-style attack). Eykholt et al. [14] showed sticker-based perturbations that reliably fool traffic-sign recognition in the real world.
@@
While most exciting literature focuses on developing efficient algorithms to solve adversarial learning model, understanding the mechanism is also extremely crucial in order to improve the robustness of learning models. Below we briefly review some related work that motivates our work in this paper. Moosavi-Dezfooli et al. [15] introduced the DeepFool attack, viewing adversarial perturbations as minimal crossings of local decision boundaries. By iteratively linearizing the classifier, it linked perturbations to decision region geometry and provided a principled way to approximate the smallest such perturbations. Carlini and Wagner [16] formalized perturbation attacks as optimization problems with tailored loss functions, showing attack strength depends critically on objective function and constraint choices. Ilyas et al. [17] offered a feature-based view, arguing that adversarial perturbations exploit non-robust features: predictive yet human-imperceptible statistical patterns. This shifted the focus from geometry to data representation. Su, Li and Cui [18] systematically studied three types of adversarial perturbations, deriving the explicit solutions for sample-adversarial perturbations (sAP), class-universal adversarial perturbations (cuAP) and universal adversarial perturbations (uAP) for binary classification, and approximate the solution for uAP multi-classification case. Later Su and Li [19] addressed the difficulty of generating sAP for nonlinear SVMs via implicit mapping by transforming the perturbation optimization into a solvable nonlinear KKT system.
@@
On the other hand, clustering is a popular yet basic model in machine learning. Various approaches have been proposed to solve clustering problems, including K-means [20, 21], K-medoids [22, 23], hierarchical clustering [24, 25], convex clustering [26, 27, 28, 29, 30] and so on. Among them, convex clustering model [31] attracts great interest due to several advantages, such as the uniqueness of solution and theoretical guarantee of cluster recovery. To deal with high dimensional data clustering, Yuan et al. [32] proposed a dimension reduction technique for structured sparse optimization problems. Ma et al. [33] proposed an improved robust sparse convex clustering (RSCC) model, which incorporates a novel norm-based feature normalization technique to effectively identify and eliminate outlier features. Angelidakis et al. [34] developed improved algorithms for stable instances of clustering problems with center-based objectives, including K-means, K-median and K-center. However, the adversarial attack case for clustering remains untouched.
To summarize, due to the complicated structure of different learning models, most adversarial attacks are difficult to interpret. Therefore, a natural question is whether we can understand the attack for a simple learning model. This motivates our work. In this paper, we study the mechanism of adversarial attack on clustering models. The contribution of the paper can be summarized in three folds. Firstly, we start with the learning model and address the perturbation of learning models. We measure the changes in solution set of learning models under perturbation in terms of calmness. Taking convex clustering as an example, we show that attack can be viewed as a special way of noise which is relatively large. By a 2-way clustering example in the one-dimensional case, we observe that, for clustering problem with noised data, if the noise is relatively small, the clustering result will remain the same. When the noise is relatively large, the clustering result will change, meaning that attack happens. Secondly, we propose the unified bilevel optimization framework for adversarial learning, under which, the adversarial learning can be viewed as seeking the solutions of bilevel models. Finally, we discuss the properties of the deviation function, which is used to measure the effect of adversarial attack. We show the well-definedness of the so-called $\delta$-measure function, and illustrate by different 2-way and 3-way clustering examples.
@@
The organization of the paper is as follows. In Section 2, we investigate the perturbation of learning models and relate the sensitivity of solution set to the so-called calmness property in the context of perturbation analysis. In Section 3, we study the effect of perturbation on the convex clustering model and provide some examples. In Section 4, we propose the unified bilevel optimization model for adversarial learning. In Section 5, we study the so-called $\delta$-measure function to show the well-definedness as the measure of adversarial attack. Final conclusions are given in Section 6.
@@ 2.1 Learning Model
where $L(X,Y)$ is the objective of the training model, $F(X)\subseteq\mathcal{Y}$ is the feasible set of $Y$, which may be affected by the training data $X$. Let $Y^{*}$ be the optimal solution of (P), which may not be unique. The solution set of (P) is denoted as $S$.
@@
Support Vector Machine (SVM) for binary classification [35], where $X=\left[\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix},\cdots,\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right]\in\mathbb{R}^{(d+1)\times n},\ Y=(\omega,b)\in\mathbb{R}^{d+1}$, with the learning model
@@
Convex Clustering [31], where $X=[x_{1},\cdots,x_{n}]\in\mathbb{R}^{d\times n}$, $Y=[y_{1},\cdots,y_{n}]\in\mathbb{R}^{d\times n}$ with the learning model
@@
Remark. Note that for most complicated learning models such as convolutional neural network (CNN), it is usually difficult to write the explicit form of $L(X,Y)$ as well as the decision function $D_{Y^{*}}$. It then brings the challenge in solving the adversarial learning model and interpreting the role of adversarial attack.
@@ 2.2 Perturbation of Learning Models
Having introduced the learning model, we are ready to consider the perturbation of the learning model. For general optimization problems, the perturbation analysis is fully addressed in [36]. In the case of noise, let $X(\varepsilon)$ be the noised training data set where $\varepsilon\in\mathcal{X}$ is the perturbation (or noise), the output of the learning model under the perturbation $\varepsilon$ is as follows
@@
where $Y^{*}(\varepsilon)$ denotes the optimal solution of the learning model under perturbation $\varepsilon$ and $S(\varepsilon)$ denotes the solution set of (Pε).
It is obvious that $X(0)=X$ and $S(0)=S$. Here we would like to highlight that due to different learning models, the form of $\varepsilon$ could be different. If $\mathcal{X}=\mathbb{R}^{k\times n}$, then $\varepsilon$ could be the additive noise or multiplicative noise. However, if $\mathcal{X}$ is in the graph space in graph clustering, i.e., $\mathcal{X}=\left\{\mathcal{G}=(V,E)\ |\ |V|=n,\ E\text{ is any set of edges on }V\right\}$ , then $X(\varepsilon)$ could be the graph that removing vertices or changing set of edges from the current graph $X$.
@@
[37, Definition 2] Let $S(\varepsilon)=\arg\min_{Y\in F(X(\varepsilon))}L(X(\varepsilon),Y)$ be a multifunction with a closed graph, denoted as $\operatorname{gph}S(\varepsilon)$, and $(\bar{\varepsilon},\overline{Y})\in\operatorname{gph}S(\varepsilon)$. We say that $S(\cdot)$ is calm at $(\bar{\varepsilon},\overline{Y})$ provided that there exist neighborhoods $\mathcal{N}_{\varepsilon}$ of $\bar{\varepsilon}$ and $\mathcal{N}_{Y}$ of $\overline{Y}$, and a modulus $\mathcal{L}\geq 0$, such that
$S(\varepsilon)\cap\mathcal{N}_{Y}\subset S(\bar{\varepsilon})+\mathcal{L}\|\varepsilon-\bar{\varepsilon}\|\\ \mathbb{B}\quad\text{for all }\varepsilon\in\mathcal{N}_{\varepsilon},$
where $\mathbb{B}$ denotes the closed unit ball in $\mathcal{X}$ and $\operatorname{gph}S(\varepsilon):=\{(\varepsilon,Y):Y\in S(\varepsilon)\}$.
Based on the definition of calmness, one can see that the calmness of $S(\cdot)$ at $\varepsilon=0$ is particularly useful for measuring the changes of $S(\varepsilon)$ relative to $S(0)$. We formally give it below.
@@
$S(\cdot)$ is calm at $(0,\overline{Y})$ if there exist neighborhoods $\mathcal{N}_{\varepsilon}$ of 0, $\mathcal{N}_{Y}$ of $\overline{Y}$, and a modulus $\mathcal{L}_{0}>0$ such that
$S(\varepsilon)\cap\mathcal{N}_{Y}\subset S+\mathcal{L}_{0}\|\varepsilon\|\ \mathbb{B}\quad\text{for all }\varepsilon\in\mathcal{N}_{\varepsilon}.$
Therefore, one can see that if $S(\cdot)$ is calm at 0, then the changes in the solution set $S(\varepsilon)$ can be controlled by the changes in $\|\varepsilon\|$ (up to the scalar $\mathcal{L}_{0}$). In other words, the robustness of the learning model $L(\cdot,\cdot)$ is closely related to the calmness of the solution set $S(\cdot)$. For a set $S(\cdot)$, there are various ways to check calmness [38, 39, 40]; moreover, Zhou and So [41] provided an equivalent characterization by showing that the error bound property holds if and only if a suitably defined solution mapping is calm, and we will not discuss the details here.
Having successfully measured the changes of $S(\varepsilon)$ due to the perturbation $\varepsilon$, we move on to see whether there is any change in the decision function $D_{Y^{*}}(\cdot)$. Even when the solution set $S(\varepsilon)$ changes, it is still possible that the decision function remains the same. The reason is as follows. In many learning tasks, the decision function is discontinuous. For example, binary classification as shown in Example 1. In other words, for such situation, the perturbation $\varepsilon$ does not have effect on the learning result. We can regard $\varepsilon$ as a neglectable noise in this case. Therefore, the question we would like to ask is as follows: under what condition on $\varepsilon$, the decision function is not changed, i.e., $D_{Y^{*}(\varepsilon)}=D_{Y^{*}}$? The question is not easy to answer. As we mentioned before, this is also highly related to the specific form of the learning model. We will address this question in Section 3, by looking at the convex clustering model as an example.

### v2

# Bilevel Models for Adversarial Learning and A Case Study

@@ Abstract
Adversarial learning has been attracting more and more attention thanks to the fast development of machine learning and artificial intelligence. However, due to the complicated structure of most machine learning models, the mechanism of adversarial attacks is not well interpreted. How to measure the effect of attacks is still not quite clear. In this paper, we investigate the adversarial learning from the perturbation analysis point of view. We characterize the robustness of learning models through the calmness of the solution mapping. In the case of convex clustering models, we identify the conditions under which the clustering results remain the same under perturbations. When the noise level is large, it leads to an attack. Therefore, we propose two bilevel models for adversarial learning where the effect of adversarial learning is measured by some deviation function. Specifically, we systematically study the so-called $\delta$-measure and show that under certain conditions, it can be used as a deviation function in adversarial learning for convex clustering models. Finally, we conduct numerical tests to verify the above theoretical results as well as the efficiency of the two proposed bilevel models.
@@ 1 Introduction
With the fast development of machine learning and artificial intelligence, adversarial learning is receiving growing attention. The first striking demonstration came from Szegedy et al. [szegedy2014], who showed that imperceptible perturbations can reliably force modern neural networks to misclassify inputs, thereby exposing a fundamental vulnerability of high-capacity models. Goodfellow et al. [goodfellow2015] provided a concise explanation and practical attack algorithm so-called the Fast Gradient Sign Method (FGSM) for fast, effective adversarial example synthesis, and thus enabled a set of adversarial-training defenses. Kurakin et al. [kurakin2017] showed that adversarial training can be implemented on larger-scale datasets such as ImageNet, and further revealed that this approach leads to a significant improvement in the robustness of one-step methods. Roberts and Smyth [roberts2022byzantine] analyzed stochastic gradient descent under Byzantine adversaries, highlighting robustness issues that also arise in distributed or large-scale settings. Chhabra et al. [chhabra2020suspicion] presented a black-box adversarial attack algorithm for clustering models with linearly separable clusters. Later, Chhabra et al. [chhabra2022robustness] proposed a black-box adversarial attack against deep clustering models. We refer to [yuan2019adversarial, costa2024deep, zhao2025data] for the review and monographs of adversarial learning.
@@
Following widely adopted taxonomies, adversarial attacks can be divided into three types based on attacker’s knowledge: white-box (full access to parameters/gradients), black-box (only queries or input–output pairs), and gray-box in between. Due to different goals of attacks, it can also be classified as confidence reduction, untargeted misclassification and targeted misclassification. A large body of work has been devoted to designing powerful white-box attacks such as FGSM [goodfellow2015] and its iterative variants and decision-boundary-based attacks like DeepFool [moosavi2016]. Feinman et al. [feinman2017DetectingAS] proposed to detect adversarial samples using kernel density estimates in the feature space of the last hidden layer together with Bayesian uncertainty estimates. Ahmed et al. [ahmed2024reevaluating] systematically evaluate four representative deep learning attacks against intrusion-detection models on cybersecurity datasets and compare several defense strategies, highlighting how different combinations of attacks and defenses affect the robustness and accuracy of such systems. Madry et al. [madry2018pgd] cast adversarial training as min–max robust optimization and used multi-step PGD as a strong white-box baseline for $l_{p}$-bounded threats. Dong et al. [dong2018mifgsm] added momentum to iterative gradients to improve black-box transferability over standard iterative attacks. Brendel et al. [brendel2018decision] proposed boundary attack that needs only top-1 decisions, walking along the boundary while shrinking perturbations. Papernot et al. [papernot2016jsma] targeted a specific label by perturbing a small, saliency chosen set of pixels (sparse $L_{0}$-style attack). Eykholt et al. [eykholt2018robust] showed sticker-based perturbations that reliably fool traffic-sign recognition in the real world.
@@
While most exciting literature focuses on developing efficient algorithms to solve adversarial learning model, understanding the mechanism is also extremely crucial in order to improve the robustness of learning models. Below we briefly review some related literature that motivates our work in this paper. Moosavi-Dezfooli et al. [moosavi2016] introduced the DeepFool attack, viewing adversarial perturbations as minimal crossings of local decision boundaries. By iteratively linearizing the classifier, it linked perturbations to decision region geometry and provided a principled way to approximate the smallest such perturbations. Carlini and Wagner [carlini2017] formalized perturbation attacks as optimization problems with tailored loss functions, showing attack strength depends critically on objective function and constraint choices. Ilyas et al. [ilyas2019] offered a feature-based view, arguing that adversarial perturbations exploit non-robust features: predictive yet human-imperceptible statistical patterns. This shifted the focus from geometry to data representation. Su, Li and Cui [su2022optimization] systematically studied three types of adversarial perturbations, deriving the explicit solutions for sample-adversarial perturbations (sAP), class-universal adversarial perturbations (cuAP) and universal adversarial perturbations (uAP) for binary classification, and approximate the solution for uAP multi-classification case. Later Su and Li [su2022efficient] addressed the difficulty of generating sAP for nonlinear SVMs via implicit mapping by transforming the perturbation optimization into a solvable nonlinear KKT system.
On the other hand, clustering is a popular yet basic model in machine learning. The study of clustering technology dates back at least to the pioneering work of Driver and Kroeber [driver1932quantitative], who quantified cultural relationships by similarity analysis in anthropology, as well as Tryon’s monograph on cluster analysis in psychology [tryon1939cluster]. Since then, many clustering algorithms have been developed, including early agglomerative hierarchical methods based on nearest-neighbour graphs and minimum-variance criteria [florek1951liaison, ward1963hierarchical], the K-means algorithm and its modern variants [macqueen1967some, lloyd1982least], density-based methods such as DBSCAN [ester1996dbscan] and OPTICS [ankerst1999optics], and the more recent density peaks clustering algorithm (DPC) [rodriguez2014dpc, wei2023dpc]. We refer the reader to the surveys and monographs [xu2005survey, everitt2011cluster] for the vast literature on clustering methods and their applications. Below we mainly focus on convex clustering, which is most related to our paper.
@@
Convex clustering model was initially proposed by Pelckmans et al. [pelckmans2005convex], and was studied in [lindsten2011clustering, hocking2011clusterpath]. The idea of convex clustering model is as follows. If two observations belong to the same cluster, then their corresponding centroids should be the same. Convex clustering model [hocking2011clusterpath, lindsten2011clustering, chi2015splitting, sun2025resistant, panahi2017clustering, sun2021convex] has several advantages, such as the uniqueness of solution and theoretical guarantee of cluster recovery. To deal with high dimensional data clustering, Yuan et al. [yuan2022dimension] proposed a dimension reduction technique for structured sparse optimization problems. Ma et al. [ma2023improved] proposed an improved robust sparse convex clustering (RSCC) model, which incorporates a novel norm-based feature normalization technique to effectively identify and eliminate outlier features. Angelidakis et al. [angelidakis2017algorithms] developed improved algorithms for stable instances of clustering problems with center-based objectives, including K-means, K-median and K-center. However, the adversarial attack case for clustering remains untouched.
To summarize, due to the complicated structure of different learning models, most adversarial attacks are difficult to interpret. Therefore, a natural question is whether we can understand the attack for a simple learning model. This motivates our work. In this paper, we study the mechanism of adversarial attack on clustering models. The contribution of the paper can be summarized as follows. (i) Firstly, we start with the perturbation of learning models, and we qualify the robustness of the solution set of learning models under perturbation by calmness. By doing so, the robustness of learning models can be clearly analyzed by the calmness property of the solution mapping. (ii) As a case study, we identify the conditions for the robustness of convex clustering model. That is, under those conditions, the convex clustering model provides unchanged clustering results under perturbation. (iii) We introduce two bilevel optimization models for adversarial learning. In particular, we formulate two bilevel adversarial learning models based on the convex clustering model. (iv) We propose and analyze deviation measures that quantify the impact of adversarial perturbations. In particular, we study the so-called $\delta$-measure and examine its behaviour in a series of 2-way and 3-way clustering examples. We also generalize it to arbitrary K-way clustering. (v) Finally, we verify the above theoretical results and the efficiency of two proposed models by numerical results on the state-of-the-art datasets. The experiments show that convex clustering is robust to moderate perturbations, while larger perturbations produce a clear staircase behaviour of deviation $\delta$ and $N_{\mathrm{chg}}$. We also compare the direct method and fmincon on the different UCI datasets and find that RI-based function can be used as a reasonable deviation function, whereas NMI can’t.
@@
The organization of the paper is as follows. In Section 2, we investigate the perturbation of learning models and relate the sensitivity of solution mapping to the so-called calmness property in the context of perturbation analysis. In Section 3, we study the effect of perturbation on the convex clustering model and provide some examples. In Section 4, we propose the bilevel optimization models for adversarial learning. In Section 5, we study the so-called $\delta$-measure function to show the well-definedness as the measure of adversarial attack. In Section 6, we conduct numerical experiments to verify the theoretical results and the efficiency of bilevel models. Final conclusions are given in Section 7.
@@ 2.1 Learning Model
where $L(X,Y)$ is the objective of the training model, $F(X)\subseteq\mathcal{Y}$ is the feasible set of $Y$, which may be affected by the training data $X$. Let $Y^{*}$ be the optimal solution of (P), which may not be unique. The solution set of (P) is denoted as $\mathcal{S}$.
@@
Support Vector Machine (SVM) for binary classification [cortes1995support], where $X=\left[\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix},\cdots,\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right]\in\mathbb{R}^{(d+1)\times n},\ Y=(\omega,b)\in\mathbb{R}^{d+1}$, with the learning model
@@
Convex Clustering [sun2021convex], where $X=[x_{1},\cdots,x_{n}]\in\mathbb{R}^{d\times n}$, $Y=[y_{1},\cdots,y_{n}]\in\mathbb{R}^{d\times n}$ with the learning model
@@
Remark 1.
For most complicated learning models such as convolutional neural network (CNN), it is usually difficult to write down an explicit objective function $L(X,Y)$ and the associated decision function $D_{Y^{*}}$. This makes it challenging both to solve the corresponding bilevel adversarial learning models and to interpret the role of adversarial perturbations in a precise way.
@@ 2.2 Perturbation of Learning Models
Having introduced the learning model, we are ready to consider the perturbation of the learning model. For general optimization problems, the perturbation analysis is fully addressed in [bonnans2013perturbation]. In the case of noise, let $X(\varepsilon)$ be the noised training data set where $\varepsilon\in\mathcal{X}$ is the perturbation (or noise), the output of the learning model under the perturbation $\varepsilon$ is as follows.
@@
where $F(X(\varepsilon))\subseteq\mathcal{Y}$ is the feasible set of $Y$, which may be affected by the noised training data $X(\varepsilon)$. Let $Y^{*}(\varepsilon)$ denotes the optimal solution of the learning model under perturbation $\varepsilon$ and $\mathcal{S}(\varepsilon)$ denotes the solution set of (Pε).
It is obvious that $X(0)=X$ and $S(0)=S$. Here we would like to highlight that due to different learning models, the form of $\varepsilon$ could be different. If $\mathcal{X}=\mathbb{R}^{k\times n}$, then $\varepsilon$ could be the additive noise or multiplicative noise. The noise can also be point-wise modifications or feature-level modifications. If $\mathcal{X}$ is in the graph space in graph clustering, i.e., $\mathcal{X}=\left\{\mathcal{G}=(V_{\mathcal{G}},E)\ |\ |V_{\mathcal{G}}|=n,\ E\text{ is any set of edges on }V_{\mathcal{G}}\right\}$ , then $X(\varepsilon)$ could be the graph-based perturbation that removing vertices or changing some edges from the current graph $X$.
@@
[flegel2007optimality, Definition 2] Let $\mathcal{S}(\varepsilon)=\arg\min_{Y\in F(X(\varepsilon))}L(X(\varepsilon),Y)$ be a multifunction with a closed graph, denoted as $\operatorname{gph}\mathcal{S}(\varepsilon)$, and $(\bar{\varepsilon},\overline{Y})\in\operatorname{gph}\mathcal{S}(\varepsilon)$. We say that $\mathcal{S}(\cdot)$ is calm at $(\bar{\varepsilon},\overline{Y})$ provided that there exist neighborhoods $\mathcal{N}_{\varepsilon}$ of $\bar{\varepsilon}$ and $\mathcal{N}_{Y}$ of $\overline{Y}$, and a modulus $\mathcal{L}\geq 0$, such that
$\mathcal{S}(\varepsilon)\cap\mathcal{N}_{Y}\subset\mathcal{S}(\bar{\varepsilon})+\mathcal{L}\|\varepsilon-\bar{\varepsilon}\|\\ \mathbb{B}\quad\text{for all }\varepsilon\in\mathcal{N}_{\varepsilon},$
where $\mathbb{B}$ denotes the closed unit ball in $\mathcal{X}$ and $\operatorname{gph}\mathcal{S}(\varepsilon):=\{(\varepsilon,Y):Y\in S(\varepsilon)\}$.
Based on the definition of calmness, one can see that the calmness of $\mathcal{S}(\cdot)$ at $\varepsilon=0$ is particularly useful for measuring the changes of $\mathcal{S}(\varepsilon)$ relative to $\mathcal{S}(0)$. We formally give it below.
@@
$\mathcal{S}(\cdot)$ is calm at $(0,\overline{Y})$ if there exist neighborhoods $\mathcal{N}_{\varepsilon}$ of 0, $\mathcal{N}_{Y}$ of $\overline{Y}$, and a modulus $\mathcal{L}_{0}>0$ such that
$\mathcal{S}(\varepsilon)\cap\mathcal{N}_{Y}\subset\mathcal{S}+\mathcal{L}_{0}\|\varepsilon\|\ \mathbb{B}\quad\text{for all }\varepsilon\in\mathcal{N}_{\varepsilon}.$
Therefore, one can see that if $\mathcal{S}(\cdot)$ is calm at 0, then the changes in the solution set $\mathcal{S}(\varepsilon)$ can be controlled by the changes in $\|\varepsilon\|$ (up to the scalar $\mathcal{L}_{0}$). In other words, the robustness of the learning model $L(\cdot,\cdot)$ is closely related to the calmness of the solution set $\mathcal{S}(\cdot)$. For a set $\mathcal{S}(\cdot)$, there are various ways to check calmness [gfrerer2011subregcalm, kruger2015holderSubreg, gfrerer2016implicit]; moreover, Zhou and So [zhou2017unified] provided an equivalent characterization by showing that the error bound property holds if and only if a suitably defined solution mapping is calm, and we will not discuss the details here.
Having successfully measured the changes of $\mathcal{S}(\varepsilon)$ due to the perturbation $\varepsilon$, we move on to see whether there is any change in the decision function $D_{Y^{*}}(\cdot)$. Even when the solution set $\mathcal{S}(\varepsilon)$ changes, it is still possible that the decision function remains the same. The reason is as follows. In many learning tasks, the decision function is discontinuous. For example, binary classification as shown in Example 1. In other words, for such situation, the perturbation $\varepsilon$ does not have effect on the learning result. We can regard $\varepsilon$ as a neglectable noise in this case. Therefore, the question we would like to ask is as follows: under what condition on $\varepsilon$, the decision function is not changed, i.e., $D_{Y^{*}(\varepsilon)}=D_{Y^{*}}$? The question is not easy to answer. As we mentioned before, this is also highly related to the specific form of the learning model. We will address this question in Section 3, by looking at the convex clustering model as an example.